OpenStudy (anonymous):

College Algebra: I'm doing 3X3 Linear Systems and I'm having trouble solving this one. It's like everything I'm doing isn't working. The equation is: 2x+y-2z=-15 4x-2y+z=15 x+3y+2z=-5

jimthompson5910 (jim_thompson5910):

notice how the last equation has x+3y+2z=-5 where the coefficient on the x is 1 this allows us to easily isolate x without leading to fractions Let's solve x+3y+2z=-5 for x x+3y+2z=-5 x+3y+2z-3y=-5-3y x + 2z = -5 - 3y x + 2z - 2z = -5 - 3y - 2z x = -5 - 3y - 2z making sense so far?

OpenStudy (anonymous):

Alright, I understand that so far

jimthompson5910 (jim_thompson5910):

now you take that and plug it into the other equations plug x = -5 - 3y - 2z into the first equation 2x+y-2z=-15 2(-5 - 3y - 2z)+y-2z=-15 ... replace x with -5 - 3y - 2z -10-6y-4z+y-2z = -15 -5y-6z-10 = -15 -5y-6z-10+10 = -15+10 -5y-6z = -5

jimthompson5910 (jim_thompson5910):

so 2x+y-2z=-15 turns into -5y-6z = -5 after you plug in x = -5 - 3y - 2z and simplify

jimthompson5910 (jim_thompson5910):

what happens when you do the same for 4x-2y+z=15 ?

OpenStudy (anonymous):

Give me a second

jimthompson5910 (jim_thompson5910):

Sure thing. Take your time.

OpenStudy (anonymous):

So I plug x=-5-3y-2z into the second equation?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

replace all of x with -5 - 3y - 2z

OpenStudy (anonymous):

I got -14y-7z=35

jimthompson5910 (jim_thompson5910):

me too

jimthompson5910 (jim_thompson5910):

so you now have this simplified 2x2 system -5y-6z = -5 -14y-7z=35 do you see how to solve that?

OpenStudy (anonymous):

The way I see it; I need to make one of the equation be y=something or z=something. Then plug that into one of the equations

OpenStudy (anonymous):

I don't see how without getting a fraction

jimthompson5910 (jim_thompson5910):

yes, essentially you are eliminating variables to try to get an equation of 1 variable so you can solve for that

jimthompson5910 (jim_thompson5910):

which method do you want to use? elimination? or substitution?

OpenStudy (anonymous):

So doing that, I'll get 3 answers? Honestly, I don't see how I'm going to get 3 solutions from doing this

jimthompson5910 (jim_thompson5910):

you should get one solution per variable (x, y, z)

jimthompson5910 (jim_thompson5910):

were you able to solve -5y-6z = -5 -14y-7z=35 ?

OpenStudy (anonymous):

Yeah that's what I meant. So, what do I do next? I'm not sure how to solve this.

jimthompson5910 (jim_thompson5910):

let's use elimination. That may be easier

jimthompson5910 (jim_thompson5910):

which variable do you want to eliminate? y or z?

OpenStudy (anonymous):

I'll say z

jimthompson5910 (jim_thompson5910):

ok to eliminate z, we need something like -6z and +6z so they add to 0z and go away but we have -6z and -7z, so they don't add to 0z like we want them to

jimthompson5910 (jim_thompson5910):

but we can fix that. we can multiply the -6z by 7 to get -42z then we can multiply the -7z by -6 to get +42z now they finally add to 0z

jimthompson5910 (jim_thompson5910):

see how I did that?

OpenStudy (anonymous):

Wait, yes I do.

jimthompson5910 (jim_thompson5910):